OpenStudy (anonymous):

integrate t^2. e((t^3)/3)

OpenStudy (anonymous):

\[\large\int t^2~e^{\frac{t^3}{3}}~dt\] Let \(u=\dfrac{t^3}{3}~\Rightarrow~du=t^2~dt\) \[\large\int e^u~du\]